OpenStudy (anonymous):

Simplify sqrt(27) - sqrt(3) + 2[sqrt(12)]

OpenStudy (anonymous):

sqrt(27)-sqrt(3)+2sqrt(12) 3sqrt(3)-sqrt(3)+4sqrt(3) 6sqrt

OpenStudy (anonymous):

6sqrt(3)

OpenStudy (anonymous):

\[\sqrt{3*9}-\sqrt{3}+2\sqrt{4*3}=3\sqrt{3}-\sqrt{3}+4\sqrt{3}=6\sqrt{3}\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Simplify [sqrt(36) - 4][sqrt(36) + 4]

OpenStudy (anonymous):

is the 4 inside the radical?

OpenStudy (anonymous):

If not then you will foil this [sqrt(36)-4][sqrt(36)+4] 36-4sqrt(36)+4sqrt(36)-16, the -4sqrt(36) adn 4sqrt(36) cancel 36-16 20

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that +20

OpenStudy (anonymous):

Yes it is